OpenStudy (anonymous):

i got this equation and i am a bit confused on how to get the answer... help? if x is the solution to (3x-1/3x+1)-2=0 then x^2+x=

OpenStudy (anonymous):

3x-1 /3x+1 = 2 3x-1= 2 (3x+1) 3x-1 = 6x +2 -3x =3 x=-1 so x^2 + x = (-1)^2 +(-1) = 1-1 =0

OpenStudy (anonymous):

Thank you! GOod answer !

OpenStudy (anonymous):

your welcome